OpenStudy (anonymous):

Archie has to go to school this morning for an important test, but he woke up late. He can either take the bus or take his unreliable car. If he takes the car, Archie knows from experience that he will make it to school without breaking down with 0.7 probability . However, the bus to school runs late 55% of the time. Archie decides to choose between these options by tossing a coin. Suppose that Archie does, in fact, make it to the test on time. What is the probability that he took the bus? Round your answer to two decimal places.

OpenStudy (anonymous):

\[\frac{1}{2}\times .7+\frac{1}{2}\times .45\]